# Eventually LIL Regret: Almost Sure $\ln\ln T$ Regret for a sub-Gaussian Mixture on Unbounded Data

### v2

@@ 4 Main Result: Iterated Logarithmic Conditional Regret
$\forall t\geq 1,\quad R_{t}\leq\frac{S_{t}^{2}}{2V_{t}}\leq C\left(1+\frac{1+\ln^{2}(1/\alpha)}{V_{t}}+\ln(1/\alpha)+\ln\ln\left(c\sqrt{1+V_{t}}\right)\right).$
@@
We now show that one can even arrive at conditional regret bounds similar to those in (4.1), starting from CS for mean. To this end, consider again the special case where the data is drawn from a distribution $P$ such that $(S_{t})_{t\geq 1}$ is a sub-Gaussian process with variance proxy $(V_{t})_{t\geq 1}$. Moreover, for some predictable $\mu_{t}$, let $S_{t}=\sum_{i\leq t}(X_{i}-\mu_{i})$, and suppose we have the following CS with $\bar{\mu}_{t}=\sum_{i\leq t}\mu_{i}/t$, $\hat{\mu}_{t}=\sum_{i\leq t}X_{i}/t$, and positive constants $c_{1}$ and $c_{2}$:
$P\left[\forall t\geq 1,~\left|{\bar{\mu}_{t}-\hat{\mu}_{t}}\right|\leq\sqrt{\frac{2V_{t}\left(c_{1}\ln\ln V_{t}+c_{2}\ln\frac{1}{\alpha}\right)}{t^{2}}}\right]\geq 1-\alpha.$
Let $\bar{\mathcal{E}}_{\alpha}$ be the set where (4) holds. Recall that one approach for arriving at inequalities of the form (4) is to threshold a non-negative (mixture) super-martingale at a level $\tfrac{1}{\alpha}$, followed by an application of Ville’s inequality [Darling and Robbins, 1967a, Robbins and Siegmund, 1970]. Thus, $\bar{\mathcal{E}}_{\alpha}$ can equivalently be thought of as $\mathcal{E}_{\alpha}$ defined in (2.1). Eq. (4) can be rearranged to show that the following inequality holds path-wise on $\bar{\mathcal{E}}_{\alpha}$:
$\forall t\geq 1,\qquad\frac{S^{2}_{t}}{2V_{t}}\leq c_{1}\ln\ln V_{t}+c_{2}\ln\frac{1}{\alpha}.$
Since $R_{t}\leq\tfrac{S^{2}_{t}}{2V_{t}}$, we also get (4.1) with $C=c_{1}\vee c_{2}$. However, we point out that our direct approach in Theorem 4.1 yields tighter path-wise regret bounds that hold unconditionally (Eq. (7)), and lead to exact upper LIL, as shown later in Theorem 4.2.

### v3

@@ 4 Main Result: Iterated Logarithmic Conditional Regret
$\forall t\geq 1,\quad R_{t}\leq C\left(1+\frac{1+\ln^{2}(1/\alpha)}{V_{t}}+\ln(1/\alpha)+\ln\ln\left(c\sqrt{1+V_{t}}\right)\right).$
@@
We now explain how, in martingale-based CS constructions, one can also read off a conditional regret bound similar to those in (4.1). To this end, consider again the special case where the data is drawn from a distribution $P$ such that $(S_{t})_{t\geq 1}$ is a sub-Gaussian process with variance proxy $(V_{t})_{t\geq 1}$. Moreover, for some predictable $\mu_{t}$, let $S_{t}=\sum_{i\leq t}(X_{i}-\mu_{i})$, and suppose we have the following CS with $\bar{\mu}_{t}=\sum_{i\leq t}\mu_{i}/t$, $\hat{\mu}_{t}=\sum_{i\leq t}X_{i}/t$, a positive constant $c_{1}$, and $\alpha\in(0,1)$:
$P\left[\forall t\geq 1,~\left|{\bar{\mu}_{t}-\hat{\mu}_{t}}\right|\leq\sqrt{\frac{2V_{t}\left(c_{1}\ln\ln(e+V_{t})+\ln\frac{1}{\alpha}\right)}{t^{2}}}\right]\geq 1-\alpha.$
One frequently used approach for arriving at inequalities of the form (4) is to threshold a non-negative (mixture) super-martingale (or equivalently, a wealth process), say $W_{t}$, at a level $\tfrac{1}{\alpha}$, followed by an application of Ville’s inequality [Darling and Robbins, 1967a, Robbins and Siegmund, 1970].
As earlier, let $\mathcal{E}_{\alpha}:=\{\sup_{t}\ln W_{t}\leq\ln(1/\alpha)\}$. While directly thresholding a wealth process $W_{t}$ (as in $\mathcal{E}_{\alpha}$) would give an implicit CS for mean, to get an explicit CS of the form in (4), these martingale-based approaches typically proceed by deriving a lower bound for $\ln W_{t}$ that is relatively tight on the set $\mathcal{E}_{\alpha}$, and thresholding this lower bound process, instead. In particular, these methods can be seen to proceed by implicitly establishing an inequality like the following: on $\mathcal{E}_{\alpha}$,
$\ln W_{t}\geq\frac{S^{2}_{t}}{2V_{t}}-c_{1}\ln\ln(e+V_{t}),\quad\forall t\geq 1.$
implying
$\left\{\sup_{t}\ln W_{t}\leq\ln\frac{1}{\alpha}\right\}\subseteq\left\{\sup_{t}\left(\frac{S^{2}_{t}}{2V_{t}}-c_{1}\ln\ln(e+V_{t})\right)\leq\ln\frac{1}{\alpha}\right\},$
hence (4). But, the above inequality also implies the following on $\mathcal{E}_{\alpha}$:
$R_{t}=\frac{S^{2}_{t}}{2V_{t}}-\ln W_{t}\leq c_{1}\ln\ln(e+V_{t}),\quad\forall t\geq 1,$
yielding a conditional regret bound of $O(\ln\ln(e+V_{t}))$ on $\mathcal{E}_{\alpha}$, in the same spirit as (4.1) with $C=c_{1}$. We remark that our direct approach in Theorem 4.1 yields tighter path-wise regret bounds that hold unconditionally (Eq. (7)), and lead to exact upper LIL, as shown later in Theorem 4.2.